# RIME: Robust Preference-based Reinforcement Learning with Noisy Preferences

### v2

@@ Abstract
Preference-based Reinforcement Learning (PbRL) avoids the need for reward engineering by harnessing human preferences as the reward signal. However, current PbRL algorithms over-reliance on high-quality feedback from domain experts, which results in a lack of robustness. In this paper, we present RIME, a robust PbRL algorithm for effective reward learning from noisy preferences. Our method incorporates a sample selection-based discriminator to dynamically filter denoised preferences for robust training. To mitigate the accumulated error caused by incorrect selection, we propose to warm start the reward model, which additionally bridges the performance gap during transition from pre-training to online training in PbRL. Our experiments on robotic manipulation and locomotion tasks demonstrate that RIME significantly enhances the robustness of the current state-of-the-art PbRL method. Ablation studies further demonstrate that the warm start is crucial for both robustness and feedback-efficiency in limited-feedback cases.
@@ 1 Introduction
Reinforcement Learning (RL) has demonstrated remarkable performance in various domains, including gameplay (Vinyals et al., 2019; Perolat et al., 2022; Kaufmann et al., 2023), robotics (Chen et al., 2022), autonomous systems (Bellemare et al., 2020; Zhou et al., 2020), etc. However, the key determinant of RL success often hinges on the careful design of reward functions, which can be both labor-intensive and error-prone. In this context, Preference-Based RL (PbRL) (Akrour et al., 2011; Christiano et al., 2017) emerges as a valuable alternative, negating the need for hand-crafted reward functions. PbRL adopts a human-in-the-loop paradigm, where human teachers provide preferences over distinct agent behaviors as the reward signal.
@@
Nevertheless, existing works in PbRL have primarily focused on enhancing feedback-efficiency, aiming to maximize the expected cumulative rewards with few number of feedback queries. This focus induces a substantial reliance on high-quality feedback, typically assuming expertise on the teachers (Liu et al., 2022; Kim et al., 2022). However, humans are prone to errors (Christiano et al., 2017). In broader applications, feedback is often sourced from non-expert users or crowd-sourcing platforms, where the quality can be inconsistent and noisy. Further complicating the matter, Lee et al. (2021a) showed that even a mere 10% corruption rate in preference labels can dramatically degrade the performance. The lack of robustness to noisy preference labels hinders the wide application of PbRL.
@@
Meanwhile, learning from noisy labels, also known as robust training, is a rising concern in deep learning, as such labels severely degrade the generalization performance of deep neural networks. Song et al. (2022) classifies robust training methods into four key categories: robust architecture (Cheng et al., 2020), robust regularization (Xia et al., 2020), robust loss design (Lyu & Tsang, 2019), and sample selection (Li et al., 2020; Song et al., 2021). However, it poses challenges to incorporate these advanced methods for robust training in PbRL. This complexity arises due to the limited number of feedback, which is often restricted to only hundreds of feedback in some tasks for the sake of feedback-efficiency and cost reduction, as well as the potential distribution shift issue during RL training.
@@
Figure 1: Overview of RIME. In the pre-training phase, we warm start the reward model $\hat{r}_{\psi}$ with intrinsic rewards $r^{\text{int}}$ to facilitate a smooth transition to online training phase. Post pre-training, the policy, Q-network, and reward model $\hat{r}_{\psi}$ are all inherited as initial configurations for online training. During online training, we utilize a denoising discriminator to screen denoised preferences for robust reward learning. This discriminator employs a dynamic lower bound $\tau_{\text{lower}}$ on the KL divergence between predicted preferences $P_{\psi}$ and annotated preference labels $\tilde{y}$ to filter trustworthy samples $\mathcal{D}_{t}$, and an upper bound $\tau_{\text{upper}}$ to flip highly unreliable labels $\mathcal{D}_{f}$.
@@
In this work, we aim to improve the robustness of preference-based RL methods on noisy and quantitatively limited preferences. To this end, we present RIME: Robust preference-based reInforcement learning via warM-start dEnoising discriminator. RIME modifies the training paradigm of the reward model in widely-adopted two-phase (i.e.pre-training and online training phases) pipeline of PbRL. Figure 1 shows an overview of RIME. In particular, to empower robust learning from noisy preferences, we introduce a denoising discriminator. It utilizes dynamic lower and predefined upper bounds on the Kullback–Leibler (KL) divergence between predicted and annotated preference labels to filter samples. Further, to mitigate the accumulated error caused by incorrect filtration, we propose to warm start the reward model during the pre-training phase for a good initialization. Moreover, we find that the warm start also bridges the performance gap that occurs during the transition from pre-training to online training. Our experiments demonstrate that RIME exceeds existing baselines by a large margin in noisy conditions and considerably improves robustness.
@@
In summary, our work has three main contributions. First, we present RIME, a robust reward learning algorithm for PbRL, designed to effectively train reward models from noisy feedback—an important and realistic topic that has not been studied extensively. Second, we observe a dramatic performance gap during the transition from pre-training to online training in PbRL, and propose to warm start the reward model for seamless transition, which proves to be crucial for both robustness and feedback-efficiency in limited-feedback cases. Last, we show that RIME outperforms existing PbRL baselines under noisy feedback settings, across a diverse set of robotic manipulation tasks from Meta-World (Yu et al., 2020) and locomotion tasks from the DeepMind Control Suite (Tassa et al., 2018, 2020), and further is more suitable for non-expert humans.
@@ 2 Related work
Preference-based Reinforcement Learning. Incorporating human feedback into the training of reward models has proven effective in various domains, including natural language processing (Ouyang et al., 2022), multi-modal (Lee et al., 2023), and reinforcement learning (Christiano et al., 2017; Ibarz et al., 2018; Hejna III & Sadigh, 2023). In the context of RL, Christiano et al. (2017) proposed a comprehensive framework for PbRL. To improve feedback-efficiency, PEBBLE (Lee et al., 2021b) used unsupervised exploration for policy pre-training. SURF (Park et al., 2021) employed data augmentation and semi-supervised learning to enrich the preference dataset. RUNE (Liang et al., 2021) encouraged exploration by modulating reward uncertainty. MRN (Liu et al., 2022) introduced a bi-level optimization to optimize the Q-function’s performance. PT (Kim et al., 2022) utilized Transformer architecture to model non-Markovian rewards, proving effective in complex tasks.
@@
Despite these advancements, the focus on feedback efficiency should not overshadow the equally critical issue of robustness in PbRL. Lee et al. (2021a) indicated that a mere 10% rate of corrupted preference labels can significantly impair algorithmic performance. Moreover, in broader application scenarios, the gathering of non-expert preferences exacerbates the risk of introducing erroneous labels. Therefore, enhancing the robustness in PbRL remains a vital research direction. In this work, we address robust PbRL via a warm-start denoising discriminator, which dynamically filter denoised preferences and is more adaptable to distribution shift cases during RL training.
@@
Learning from Noisy Labels. Learning from noisy labels has gained more attention in supervised learning, particularly in light of the widespread presence of noisy or imprecise labels in real-world applications. A variety of approaches have been proposed for robust training (Song et al., 2022), including architectural modifications (Goldberger & Ben-Reuven, 2016), regularization (Lukasik et al., 2020), loss function designs (Zhang & Sabuncu, 2018), and sample selection methods (Wang et al., 2021). Despite these advancements, the direct application of these methods to reward learning in PbRL has presented challenges, notably due to the limited sample sizes and the absence of i.i.d. of samples. In PbRL, Xue et al. (2023) proposed an encoder-decoder architecture to model diverse human preferences, which required approximate 100 times amount of preference labels used in our experiments. Our approach can be situated within the sample selection category and improves robustness while preserving feedback-efficiency.
@@
Policy-to-Value Reincarnating RL. Policy-to-value reincarnating RL (PVRL) represents transferring a sub-optimal teacher policy to a value-based RL student agent (Agarwal et al., 2022). Uchendu et al. (2023) found that a randomly initialized Q-network in PVRL leads to the teacher policy being forgotten quickly. Within the widely-adopted pipeline of PbRL, the challenge intrinsic to PVRL also arise during the transition from pre-training to online training, but has been neglected in previous research (Lee et al., 2021b; Park et al., 2021; Liang et al., 2021; Liu et al., 2022). The issue of forgetting pre-training policy becomes more critical under noisy feedback conditions. Based on the observation, we propose to warm start the reward model for a seamless transition. Our ablation study demonstrates that the warm start is crucial for both robustness and feedback-efficiency.
@@ 3 Preliminaries
Preference-based Reinforcement Learning. In standard RL, an agent interacts with an environment in discrete time steps (Sutton & Barto, 2018). At each time step $t$, the agent observes the current state $\mathbf{s}_{t}$ and selects an action $\mathbf{a}_{t}$ according to its policy $\pi(\mathbf{a}_{t}|\mathbf{s}_{t})$. The environment responds by emitting a reward $r(\mathbf{s}_{t},\mathbf{a}_{t})$ and transitioning to the next state $\mathbf{s}_{t+1}$. The agent’s objective is to learn a policy that maximizes the expected return.
In Preference-based RL, there is no predefined reward function. Instead, a teacher offers preferences between agent’s behaviors and an estimated reward function $\hat{r}_{\psi}$ is trained to align with collected preferences. Following previous works (Lee et al., 2021b; Liu et al., 2022; Kim et al., 2022), we consider preferences over two trajectory segments of length $H$, where segment $\sigma=\{(\mathbf{s}_{1},\mathbf{a}_{1}),...,(\mathbf{s}_{H},\mathbf{a}_{H})\}$. Given a pair of segments $(\sigma^{0},\sigma^{1})$, a teacher provides a preference label $\tilde{y}$ from the set $\{(1,0),(0,1),(0.5,0.5)\}$. The label $\tilde{y}=(1,0)$ signifies $\sigma^{0}\succ\sigma^{1}$, $\tilde{y}=(0,1)$ signifies $\sigma^{1}\succ\sigma^{0}$, and $\tilde{y}=(0.5,0.5)$ represents an equally preferable case, where $\sigma^{i}\succ\sigma^{j}$ denotes that segment $i$ is preferred over segment $j$. Each feedback is stored in a dataset $\mathcal{D}$ as a triple $(\sigma^{0},\sigma^{1},\tilde{y})$. Following the Bradley-Terry model (Bradley & Terry, 1952), the preference predicted by the estimated reward function $\hat{r}_{\psi}$ is formulated as:
@@
Unsupervised Pre-training in PbRL. Pre-training agents is important in PbRL because the initial random policy often results in uninstructive preference queries, requiring lots of queries for even elementary learning progress. Recent study addressed this issue through unsupervised exploration for policy pre-training (Lee et al., 2021b). Specifically, agents are encouraged to traverse a more expansive state space by utilizing an intrinsic reward function derived from particle-based state entropy (Singh et al., 2003). Formally, the intrinsic reward is defined as (Liu & Abbeel, 2021):
@@
where $\mathbf{s}_{t}^{k}$ is the $k$-th nearest neighbor of $\mathbf{s}_{t}$. This reward motivates the agent to explore a broader diversity of states. This exploration, in turn, leads to a varied set of agent’s behaviors, facilitating more informative preference queries.
Noisy Preferences in PbRL. We denote the annotated preference labels as $\tilde{y}$ and the ground-truth preference labels, typically sourced from expert human teachers or scripted teachers, as $y$. To simulate the noise in human annotations, Lee et al. (2021a) introduced four noisy 0-1 labeling models: Equal, Skip, Myopic, and Mistake. The “Mistake” model, in particular, proved to be significantly detrimental to performance across various environments. It hypothesizes that the preference dataset is contaminated with corrupted preferences whose annotated labels $\tilde{y}$ are $1-y$. Building upon these insights, our work starts from addressing robust reward learning under the mistake model settings because of the guidance that solutions designed for complex challenges could be effectively adapted to address simpler ones.
4 Method
In this section, we formally introduce RIME: Robust preference-based reInforcement learning via warM-start dEnoising discriminator. RIME consists two main components: 1) a denoising discriminator designed to filter out corrupted preference data while accounting for training instability and distribution shift issue, and 2) a warm start method to effectively initialize the reward model and enable a seamless transition from pre-training to online training. See Figure 1 for the overview of RIME. The full procedure of RIME is detailed in Appendix A.
@@ 4.1 Denoising Discriminator
In the presence of noisy labels, it is well motivated to distinguish between clean and corrupted samples for robust training. Existing research indicates that deep neural networks first learn generalizable patterns before overfitting to the noise in the data. Therefore, prioritizing samples associated with smaller losses as clean ones is a well-founded approach to improve robustness. Inspired by this insight, we theoretically establish a lower bound on the KL divergence between the predicted preference $P_{\psi}$ and the annotated preference $\tilde{y}$ for corrupted samples, in order to filter out large-loss corrupted samples.
@@
$D_{\mathrm{KL}}\left(P_{\psi}(x)\|\tilde{y}(x)\right)\geq-\ln\rho+\frac{\rho}{2}+\mathcal{O}(\rho^{2})$ (4)
The proof of Theorem 4.1 is presented in Appendix C. Based on Theorem 4.1, we formulate the lower bound on KL divergence threshold to filter out untrustworthy samples as $\tau_{\text{base}}=-\ln\rho+\alpha\rho$ in practical, where $\rho$ denotes the maximum cross-entropy loss on trustworthy samples observed during the last update, and $\alpha$ serves as a tunable hyperparameter with a value range in $(0,0.5]$ in theoretical.
However, the shifting state distribution complicates the robust training problem in RL, compared to deep learning contexts. To add tolerance for trustworthy samples in cases of distribution shift, we introduce an auxiliary term characterizing the uncertainty for filtration, defined as $\tau_{\text{unc}}=\beta_{t}\cdot s_{\text{KL}}$, where $\beta_{t}$ is a time-dependent parameter, and $s_{\text{KL}}$ is the standard deviation of the KL divergence. Our intuition is that the inclusion of out-of-distribution data for training is likely to induce fluctuations in the training loss. Therefore, the complete threshold equation is formulated as follows:
@@
We utilize a linear decay schedule for $\beta_{t}$ to initially allow greater tolerance for samples while becoming increasingly conservative over time, i.e.$\beta_{t}=\max(\beta_{\min},\beta_{\max}-kt)$. At each training step for the reward model, we apply the threshold in Equation (5) to identify trustworthy sample dataset $\mathcal{D}_{t}$, as described below:
$\mathcal{D}_{t}=\{(\sigma^{0},\sigma^{1},\tilde{y})\,|\,D_{\mathrm{KL}}(P_{\psi}(\sigma^{0},\sigma^{1})\|\tilde{y})<\tau_{\text{lower}}\}$ (6)
@@
Beyond improving sample utilization, the label-flipping method also bolsters the model’s predictive confidence and reduce output entropy (Grandvalet & Bengio, 2004). Following two filtering steps, the reward model is trained on the unified datasets $\mathcal{D}_{t}\cup\mathcal{D}_{f}$, using the loss function as follows:
@@
Our denoising discriminator belongs to the category of sample selection methods for robust training. However, it is different in using a dynamically adjusted threshold, augmented by a term that captures distributional shifts, making it more suitable for RL training process.
(a) $\epsilon=0.1$
(b) $\epsilon=0.15$
(c) $\epsilon=0.2$
(d) $\epsilon=0.25$
(e) $\epsilon=0.3$
Figure 2: Learning curves for robotic manipulation tasks from Meta-world, where each row represents a specific task and each column corresponds to a different error rate $\epsilon$. SAC serves as a performance upper bound, using a ground-truth reward function unavailable in PbRL settings. The corresponding number of feedback in total and per session are show in Table 1. The solid line and shaded regions respectively denote mean and standard deviation of success rate, across ten runs.

### v3

@@ Abstract
Preference-based Reinforcement Learning (PbRL) circumvents the need for reward engineering by harnessing human preferences as the reward signal. However, current PbRL methods excessively depend on high-quality feedback from domain experts, which results in a lack of robustness. In this paper, we present RIME, a robust PbRL algorithm for effective reward learning from noisy preferences. Our method utilizes a sample selection-based discriminator to dynamically filter out noise and ensure robust training. To counteract the cumulative error stemming from incorrect selection, we suggest a warm start for the reward model, which additionally bridges the performance gap during the transition from pre-training to online training in PbRL. Our experiments on robotic manipulation and locomotion tasks demonstrate that RIME significantly enhances the robustness of the state-of-the-art PbRL method. Code is available at https://github.com/CJReinforce/RIME_ICML2024.
@@ 1 Introduction
Reinforcement Learning (RL) has demonstrated remarkable performance in various domains, including gameplay (Perolat et al., 2022; Kaufmann et al., 2023), robotics (Chen et al., 2022), autonomous systems (Bellemare et al., 2020; Zhou et al., 2020), multimodal (Yue et al., 2024), etc. The success of RL frequently relies on the meticulous crafting of reward functions, a process that can be both time-consuming and susceptible to errors. In this context, Preference-Based RL (PbRL) (Akrour et al., 2011; Cheng et al., 2011; Christiano et al., 2017) emerges as a valuable alternative, eliminating the requirement for manually designed reward functions. PbRL adopts a human-in-the-loop paradigm, where human teachers provide preferences over distinct agent behaviors as the reward signal.
@@
Nevertheless, existing works in PbRL have primarily focused on enhancing feedback-efficiency, aiming to maximize the expected return with few feedback queries. This focus induces a substantial reliance on high-quality feedback, typically assuming expertise on the teachers (Liu et al., 2022; Kim et al., 2022). However, humans are prone to errors (Christiano et al., 2017). In broader applications, feedback is often sourced from non-expert users or crowd-sourcing platforms, where the quality can be inconsistent and noisy. Further complicating the matter, Lee et al. (2021a) showed that even a mere 10% corruption rate in preference labels can dramatically degrade the performance. Therefore, the lack of robustness to noisy preference labels hinders the wide applicability of PbRL.
@@
Meanwhile, learning from noisy labels, also known as robust training, is a rising concern in deep learning, since such labels severely degrade the generalization performance of deep neural networks. Song et al. (2022) classifies robust training methods into four key categories: robust architecture (Cheng et al., 2020), robust regularization (Xia et al., 2020), robust loss design (Lyu & Tsang, 2019), and sample selection (Li et al., 2020; Song et al., 2021). However, it is challenging to effectively incorporate these advanced methods for robust training in PbRL. This complexity arises from the pursuit of feedback-efficiency and cost reduction, necessitating access to a limited amount of feedback. Simultaneously, the distribution shift problem during RL training undermines the assumption of i.i.d input data, a core principle that supports robust training methods in deep learning.
@@
Figure 1: Overview of RIME. In the pre-training phase, we warm start the reward model $\hat{r}_{\psi}$ with intrinsic rewards $r^{\text{int}}$ to facilitate a smooth transition to the online training phase. Post pre-training, the policy, Q-network, and reward model $\hat{r}_{\psi}$ are all inherited as initial configurations for online training. During online training, we utilize a denoising discriminator to screen denoised preferences for robust reward learning. This discriminator employs a dynamic lower bound $\tau_{\text{lower}}$ on the KL divergence between predicted preferences $P_{\psi}$ and annotated preference labels $\tilde{y}$ to filter trustworthy samples $\mathcal{D}_{t}$, and an upper bound $\tau_{\text{upper}}$ to flip highly unreliable labels $\mathcal{D}_{f}$.
@@
In this work, we aim to improve the robustness of preference-based RL methods on noisy and quantitatively limited preferences. To this end, we present RIME: Robust preference-based reInforcement learning via warM-start dEnoising discriminator. RIME modifies the training paradigm of the reward model in the widely-adopted two-phase (i.e.,pre-training and online training phases) pipeline of PbRL. Figure 1 shows an overview of RIME. In particular, to empower robust learning from noisy preferences, we introduce a denoising discriminator. It utilizes dynamic lower and predefined upper bounds on the Kullback–Leibler (KL) divergence between predicted and annotated preference labels to filter samples. Further, to mitigate the accumulated error caused by incorrect filtration, we propose to warm start the reward model during the pre-training phase for a good initialization. Moreover, we find that the warm start also bridges the performance gap that occurs during the transition from pre-training to online training. Our experimental results indicate that RIME significantly outperforms existing baselines under noisy preference conditions, thereby substantially enhancing robustness for PbRL.
@@
In summary, our work has three main contributions. First, we present RIME, a robust reward learning algorithm for PbRL, designed to effectively train reward models from noisy feedback—an important and realistic topic that has not been studied extensively. Second, we observe a dramatic performance gap during the transition from pre-training to online training in PbRL and propose to warm start the reward model for a seamless transition, which proves to be crucial for both robustness and feedback-efficiency in limited-feedback cases. Last, we show that RIME outperforms existing PbRL baselines under noisy feedback settings, across a diverse set of robotic manipulation tasks from Meta-World (Yu et al., 2020) and locomotion tasks from the DeepMind Control Suite (Tassa et al., 2018, 2020), and further is more suitable for non-expert humans.
@@ 2 Related work
Preference-based Reinforcement Learning. Incorporating human feedback into the training of reward models has proven effective in various domains, including natural language processing (Ouyang et al., 2022), multimodal (Lee et al., 2023), and reinforcement learning (Christiano et al., 2017; Ibarz et al., 2018; Hejna III & Sadigh, 2023). In the context of RL, Christiano et al. (2017) proposed a comprehensive framework for PbRL. To improve feedback-efficiency, PEBBLE (Lee et al., 2021b) used unsupervised exploration for policy pre-training. SURF (Park et al., 2021) employed data augmentation and semi-supervised learning to enrich the preference dataset. RUNE (Liang et al., 2021) encouraged exploration by modulating reward uncertainty. MRN (Liu et al., 2022) introduced a bi-level optimization to optimize the Q-function’s performance. PT (Kim et al., 2022) utilized Transformer architecture to model non-Markovian rewards, proving effective in complex tasks.
@@
Despite these advancements, the focus on feedback efficiency should not overshadow the equally critical issue of robustness in PbRL. Lee et al. (2021a) indicated that a mere 10% rate of corrupted preferences can significantly impair algorithmic performance. Furthermore, in more extensive application contexts, the collection of preferences from non-experts increases the likelihood of incorporating incorrect labels. Therefore, enhancing the robustness of PbRL remains a vital research direction. In this work, we address robust PbRL via a warm-start denoising discriminator, which dynamically filters denoised preferences and is more adaptable to cases of distribution shift during RL training.
@@
Learning from Noisy Labels. Learning from noisy labels has gained more attention in supervised learning, particularly in light of the wide presence of noisy or imprecise labels in real-world applications. A variety of approaches have been proposed for robust training (Song et al., 2022), including architectural modifications (Goldberger & Ben-Reuven, 2016), regularization (Lukasik et al., 2020), loss function designs (Zhang & Sabuncu, 2018), and sample selection methods (Wang et al., 2021). Despite these advancements, the direct application of these methods to reward learning in PbRL has presented challenges, mainly due to the limited sample sizes and the absence of i.i.d. of samples. In the context of PbRL, Xue et al. (2023) proposed an encoder-decoder architecture to model diverse human preferences, which required approximately 100 times the amount of preference labels used in our experiments. Our approach can be situated within the sample selection category and improves robustness while preserving feedback-efficiency.
@@
Policy-to-Value Reincarnating RL. Policy-to-value reincarnating RL (PVRL) transfers a sub-optimal teacher policy to a value-based RL student agent (Agarwal et al., 2022). Uchendu et al. (2023) found that a randomly initialized Q-network in PVRL leads to the teacher policy being forgotten quickly. Within the widely-adopted pipeline of PbRL, the challenge intrinsic to PVRL also arises during the transition from pre-training to online training, but has been neglected in previous research (Lee et al., 2021b; Park et al., 2021; Liang et al., 2021; Liu et al., 2022). The issue of forgetting the pre-training policy becomes more critical under noisy feedback conditions, as detailed in Section 4.2. Based on the observation, we propose to warm start the reward model for a seamless transition. Our ablation study demonstrates that the warm start is crucial for both robustness and feedback-efficiency.
@@ 3 Preliminaries
Preference-based Reinforcement Learning. In standard RL, an agent interacts with an environment in discrete time steps (Sutton & Barto, 2018). At each time step $t$, the agent observes the current state $\mathbf{s}_{t}$ and selects an action $\mathbf{a}_{t}$ according to its policy $\pi(\cdot|\mathbf{s}_{t})$. The environment responds by emitting a reward $r(\mathbf{s}_{t},\mathbf{a}_{t})$ and transitioning to the next state $\mathbf{s}_{t+1}$. The agent’s objective is to learn a policy that maximizes the expected return, $\mathcal{R}_{0}=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, which is defined as a discounted cumulative sum of the reward with the discount factor $\gamma$.
In Preference-based RL, there is no predefined reward function. Instead, a teacher offers preferences between the agent’s behaviors, and an estimated reward function $\hat{r}_{\psi}$ is trained to align with collected preferences. Following previous works (Lee et al., 2021b; Liu et al., 2022; Kim et al., 2022), we consider preferences over two trajectory segments of length $H$, where segment $\sigma=\{(\mathbf{s}_{1},\mathbf{a}_{1}),...,(\mathbf{s}_{H},\mathbf{a}_{H})\}$. Given a pair of segments $(\sigma^{0},\sigma^{1})$, a teacher provides a preference label $\tilde{y}$ from the set $\{(1,0),(0,1),(0.5,0.5)\}$. The label $\tilde{y}=(1,0)$ signifies $\sigma^{0}\succ\sigma^{1}$, $\tilde{y}=(0,1)$ signifies $\sigma^{1}\succ\sigma^{0}$, and $\tilde{y}=(0.5,0.5)$ represents an equally preferable case, where $\sigma^{i}\succ\sigma^{j}$ denotes that segment $i$ is preferred over segment $j$. Each feedback is stored in a dataset $\mathcal{D}$ as a triple $(\sigma^{0},\sigma^{1},\tilde{y})$. Following the Bradley-Terry model (Bradley & Terry, 1952), the preference predicted by the estimated reward function $\hat{r}_{\psi}$ is formulated as:
@@
Unsupervised Pre-training in PbRL. Pre-training the agent is important in PbRL because the initial random policy often results in uninstructive preference queries, requiring many queries for even elementary learning progress. Recent studies addressed this issue through unsupervised exploration for policy pre-training (Lee et al., 2021b). Specifically, agents are encouraged to traverse a more expansive state space by using an intrinsic reward derived from particle-based state entropy (Singh et al., 2003). Formally, the intrinsic reward is defined as (Liu & Abbeel, 2021):
@@
where $\mathbf{s}_{t}^{k}$ is the $k$-th nearest neighbor of $\mathbf{s}_{t}$. This reward motivates the agent to explore a broader diversity of states. This exploration, in turn, leads to a varied set of agent behaviors, facilitating more informative preference queries.
Noisy Preferences in PbRL. We denote the annotated preference labels as $\tilde{y}$ and the ground-truth preference labels, typically sourced from expert human teachers or scripted teachers, as $y$. To simulate the noise in human annotations, Lee et al. (2021a) introduced four noisy 0-1 labeling models: Equal, Skip, Myopic, and Mistake. The “Mistake” model, in particular, proved to be significantly detrimental to performance across various environments. It hypothesizes that the preference dataset is contaminated with corrupted preferences whose annotated labels $\tilde{y}$ are $1-y$. Drawing on previous insights, our work starts from addressing robust reward learning under the “Mistake” model settings. This approach is guided by empirical evidence suggesting that solutions developed to overcome complex challenges could be efficiently adapted to simpler cases.
4 RIME
In this section, we formally introduce RIME: Robust preference-based reInforcement learning via warM-start dEnoising discriminator. RIME consists of two main components: 1) a denoising discriminator designed to filter out corrupted preferences while accounting for training instability and distribution shifts, and 2) a warm start method to effectively initialize the reward model and enable a seamless transition from pre-training to online training. See Figure 1 for the overview of RIME. The full procedure of RIME is detailed in Appendix A.
@@ 4.1 Denoising Discriminator
In the presence of noisy labels, it is well-motivated to distinguish between clean and corrupted samples for robust training. Existing research indicates that deep neural networks first learn generalizable patterns before overfitting to the noise in the data (Arpit et al., 2017; Li et al., 2020). Therefore, prioritizing samples associated with smaller losses as clean ones is a well-founded approach to improve robustness. Inspired by this, a theoretical lower bound on the KL divergence between the predicted preference $P_{\psi}$ and the annotated preference $\tilde{y}$ for corrupted samples could be established to filter out large-loss corrupted samples.
@@
$D_{\mathrm{KL}}\left(\tilde{y}(x)\|P_{\psi}(x)\right)\geq-\ln\rho+\frac{\rho}{2}+\mathcal{O}(\rho^{2})$ (4)
The proof of Theorem 4 is presented in Appendix C. Based on Theorem 4, the lower bound on KL divergence threshold could be formulated to filter out untrustworthy samples as $\tau_{\text{base}}=-\ln\rho+\alpha\rho$ in practice, where $\rho$ denotes the maximum cross-entropy loss on trustworthy samples observed during the last update, and $\alpha$ is a tunable hyperparameter with a theoretically-determined value range in $(0,0.5]$.
However, compared to deep learning, the shifting state distribution makes the robust training problem in RL more complicated. To add tolerance for clean samples in cases of distribution shift, we introduce an auxiliary term characterizing the uncertainty for filtration, defined as $\tau_{\text{unc}}=\beta_{t}\cdot s_{\text{KL}}$, where $\beta_{t}$ is a time-dependent parameter, and $s_{\text{KL}}$ is the standard deviation of the KL divergence. Our intuition is that the inclusion of out-of-distribution data for training is likely to induce fluctuations in training loss. Therefore, the complete threshold equation is formulated as:
@@
We utilize a linear decay schedule for $\beta_{t}$ to initially allow greater tolerance for samples while becoming increasingly conservative over time, i.e.,$\beta_{t}=\max(\beta_{\min},\beta_{\max}-kt)$. At each training step for the reward model, we apply the threshold in Equation (5) to identify trustworthy sample dataset $\mathcal{D}_{t}$, as described below:
$\mathcal{D}_{t}=\{(\sigma^{0},\sigma^{1},\tilde{y})\,|\,D_{\mathrm{KL}}(\tilde{y}\|P_{\psi}(\sigma^{0},\sigma^{1}))<\tau_{\text{lower}}\}$ (6)
@@
Beyond improving sample utilization, the label-flipping method also bolsters the model’s predictive confidence and reduces output entropy (Grandvalet & Bengio, 2004). Following two filtering steps, the reward model is trained on the unified datasets $\mathcal{D}_{t}\cup\mathcal{D}_{f}$, using the loss function as follows:
@@
Our denoising discriminator belongs to the category of sample selection methods for robust training. It stands out due to its use of a dynamically adjusted threshold enhanced by a term that accounts for instability and distributional shifts, thereby making it more suitable for the RL training process.